OpenStudy (anonymous):

81x^2-72x+16 is a perfect square trinomial? True or False

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

yeah , here we go!

OpenStudy (anonymous):

what number squared is 81?

OpenStudy (anonymous):

81=9x9 16=4x4 right

OpenStudy (anonymous):

yes right so far

OpenStudy (anonymous):

so your final job is to check to see what \(2\times 9\times 4\) is

OpenStudy (anonymous):

72

OpenStudy (anonymous):

then you win !

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

\[81x^2-72x+16=(9x-4)^2\] yw